OpenStudy (anonymous):

will fan/medal/appreciate your help! systems of linear equalities. For questions1-2, what is the graph of the system? 1. y less than or equal to x+4 2x+y less than or equal to -4 2.y less than or equal to -x - 1 y greater than or equal to 2x+4

OpenStudy (texaschic101):

y <= x + 4 y intercept = (0,4) x intercept = (-4,0) plot these points, draw a solid line, and shade under the line 2x + y <= -4 y <= -2x - 4 y intercept = (0,-4) x intercept = (-2,0) plot these points, connect them with a solid line, and shade under the line if you want to see what the graph looks like, go here, and write your inequalities and they will be graphed. https://www.mathway.com/graph ======================= y <= -x - 1 y intercept = (0,-1) x intercept = (-1,0) plot these points, connect with a solid line, shading will be below the line y >= 2x + 4 y intercept = (0,4) x intercept = (-2,0) plot these points, connect with a solid line, shading will be above the line

OpenStudy (anonymous):

thank you so much!

OpenStudy (texaschic101):

no problem :)